# Disentangling the anisotropic radio sky: Fisher forecasts for 21cm arrays

### v1

@@ 3.3 Results
We assume three different radio skies, with the respective RSB components described by RSB 1, 2 and 3, while all three cases are identical in the remaining four components. We investigate the ability of HERA to constrain these RSB model parameters in two scenarios: (1) without using any priors; (2) using priors on the four foreground components (i.e. 10% uncertainties on each foreground parameter).
Due to the numerical challenges of inverting large covariance matrices when computing Fisher matrices, we use the universal SED approximation (up to 00th order) to reduce the matrix size, as explained in Sec. 2.3.1. In this toy model analysis, we only consider the frequency structure up to the zeroth order, i.e. the power law. In the case of no prior on any parameter of any component, the result of the Fisher analysis shows that the RSB model can hardly be constrained with the experimental setup in this work, as shown in Fig. 4. On the other hand, the right column of the same figure, as well as Fig. 5, shows that when we apply the priors to the uncertainties of the foreground parameters, the constraints on the RSB parameters become tighter. In particular, RSB 1 and 3 are constrained with a sub-dominant level of uncertainty, while RSB 2 is still far from being well constrained. In order to constrain a class of models similar to RSB 2, we need to further reduce the instrumental noise or give a higher priority to the foregrounds. An alternative strategy is to introduce priors on the RSB if necessary. In theory, if the uncertainty caused by the cosmic variance of a set of SH modes does not prevent us from constraining a particular RSB model with sufficient accuracy, then experiments measuring these modes can be used if the noise is low enough.
@@ 4 Discussion and conclusions
The presence of the excess radio background, or RSB, awaits further independent confirmation beyond difficult absolute spectrometry measurements. Anisotropy measurements with 21-cm arrays are a promising way to further constrain its origin. In this work, we have developed a Fisher Forecast formalism for evaluating the potential for 21-cm interferometric arrays to disentangle the anisotropic diffuse radio sky. In particular, we predict that a HERA-like array should be able to constrain the angular power spectrum (APS) of the RSB to a high degree even in the presence of multiple foreground components, with an uncertainty that depends on the true APS of the different components (Fig. 4 shows toy model cases).
@@
for a given fiducial composition of the radio sky. Ideally, we can maximise the likelihood function without using any informative priors, essentially putting all components of the sky on an equal footing. Extracting the RSB in this way (or perhaps also with the addition of ARCADE 2-based priors on some of the components) would in principle permit us to establish the existence of the RSB in a way that is independent of existing methods and results from monopole measurements. On the other hand, this component separation strategy will also be useful in deepening our understanding and tightening the constraints on all the remaining Galactic and extragalactic components, which would also facilitate other radio cosmology surveys that need to characterise and remove foreground emission.
@@
If the existence of an RSB background is confirmed, it would significantly change the interpretation of limits reported by 21-cm surveys (e.g. Acharya et al., 2023; Fialkov & Barkana, 2019), which uses the primordial radiation background as a benchmark and relies on phenomenological foreground models for science extraction. If the potential background is emitted at several different redshifts, even before or during the Epoch of Reionization (EoR), then understanding the RSB becomes a prerequisite for understanding the 21-cm signal.
@@
Discriminating between different theoretical explanations by measuring the angular power spectrum of RSBs may require a tighter constraint and thus a more elaborate experimental setup than simply detecting the presence of RSBs. Our formalism provides two steps for uncertainty assessment to efficiently design observational strategies. The first step is to select the appropriate set of spherical harmonic modes, as well as the frequency range and frequency resolution, by estimating the intrinsic uncertainty from cosmic variance (i.e. equation 9). After determining the targeted modes and frequencies, the second step is to predict the uncertainty introduced by the experimental setup (equation 16). The advantage of radio interferometry is that since each baseline is most sensitive to the fluctuation mode at a particular scale, we can choose a particular combination of baselines to tailor the analysis. For example, if the frequency anticorrelation parameter $\xi$ for a theory is highly dependent on $\ell$, so we want to constrain this theoretical model only in a particular smaller range of $\ell$ where $\xi_{\ell}\simeq\xi^{\prime}$ is approximately a constant, then we can down-weight the response to unwanted $\ell$ modes by filtering the baseline.
@@
In the cosmic variance-limited (noise-free) case where all of the spherical harmonic modes are measured within a band between $20\leq\ell\leq 90$ (or $40\leq\ell\leq 70$; see Table 1), we found that all of the foreground component parameters could be strongly constrained except for the (faint) Extragalactic Free-Free emission (Table 2). This is in the case where there is no fiducial RSB model present, and no prior has been applied. This suggests that close-packed arrays of the kind used for 21cm surveys can usefully disentangle different anisotropic components of the radio sky based on angular and spectral clustering alone, i.e. without absolute spectrometry.
@@
We also showed predictions for a HERA-like array to measure the angular power spectrum of the RSB for different toy models (results in Sec. 3.3). If the fiducial amplitude of the RSB component is of the same order as the dominant Galactic synchrotron component or larger, its amplitude and spectral/spatial power law indices can be constrained reasonably well within a typical observing season of $\sim 40$ nights or so, even if the spectral/spatial clustering properties are very similar to the Galactic synchrotron’s. This is only the case if $\sim 10\%$ priors are applied to the non-RSB component parameters, as otherwise significant degeneracies arise. If the RSB component is fainter than this at the target frequencies, however, it becomes difficult to disentangle from the foreground components based on angular and spectral clustering properties alone, and so either substantially more observing time, or stronger priors on the other components would be needed to permit its characterisation in terms of the parameters of our assumed APS model.

### v2

@@ 3.3 Results
We assume three different radio skies, with the respective RSB excess components described by RSB excess 1, 2 and 3, while all three cases are identical in the remaining four components. We investigate the ability of HERA to constrain these RSB excess parameters in two scenarios: (1) without using any priors; (2) using priors on the four foreground components (i.e. 10% uncertainties on each foreground parameter).
Due to the numerical challenges of inverting large covariance matrices when computing Fisher matrices, we use the universal SED approximation (up to 00th order) to reduce the matrix size, as explained in Sec. 2.3.1. In this toy model analysis, we only consider the frequency structure up to the zeroth order, i.e. the power law. In the case of no prior on any parameter of any component, the result of the Fisher analysis shows that the RSB excess model can hardly be constrained with the experimental setup in this work, as shown in Fig. 4. On the other hand, the right column of the same figure, as well as Fig. 5, shows that when we apply the priors to the uncertainties of the foreground parameters, the constraints on the RSB excess parameters become tighter. In particular, RSB excess 1 and 3 are constrained with a sub-dominant level of uncertainty, while RSB excess 2 is still far from being well constrained. In order to constrain a class of models similar to RSB excess 2, we need to further reduce the instrumental noise or give a higher priority to the foregrounds. An alternative strategy is to introduce priors on the RSB excess if necessary. In theory, if the uncertainty caused by the cosmic variance of a set of SH modes does not prevent us from constraining a particular RSB excess model with sufficient accuracy, then experiments measuring these modes can be used if the noise is low enough.
@@ 4 Discussion and conclusions
The origin of the RSB excess is a mystery. Anisotropy measurements with 21cm arrays are a promising way to further constrain the origin. In this work, we have developed a Fisher Forecast formalism for evaluating the potential for 21cm interferometric arrays to disentangle the anisotropic diffuse radio sky. In particular, we predict that a HERA-like array should be able to constrain the angular power spectrum (APS) of the RSB excess to a high degree even in the presence of multiple foreground components, with an uncertainty that depends on the true APS of the different components (Fig. 4 shows toy model cases).
@@
for a given fiducial composition of the radio sky. Ideally, we can maximise the likelihood function without using any informative priors, essentially putting all components of the sky on an equal footing. Extracting the RSB excess in this way (or perhaps also with the addition of ARCADE 2-based priors on some of the components) would in principle permit us to establish the existence of the RSB excess in a way that is independent of existing methods and results from monopole measurements. On the other hand, this component separation strategy will also be useful in deepening our understanding and tightening the constraints on all the remaining Galactic and extragalactic components, which would also facilitate other radio cosmology surveys that need to characterise and remove foreground emission.
@@
If the existence of an RSB excess is confirmed, it would significantly change the interpretation of limits reported by 21cm surveys (e.g. Acharya et al., 2023; Fialkov & Barkana, 2019), which uses the primordial radiation background as a benchmark and relies on phenomenological foreground models for science extraction. If the potential background is emitted at several different redshifts, even before or during the Epoch of Reionization (EoR), then understanding the RSB becomes a prerequisite for understanding the 21cm signal.
@@
Discriminating between different theoretical explanations by measuring the angular power spectrum of the RSB excess may require a tighter constraint and thus a more elaborate experimental setup than simply detecting the presence of the RSB excess. Our formalism provides two steps for uncertainty assessment to efficiently design observational strategies. The first step is to select the appropriate set of spherical harmonic modes, as well as the frequency range and frequency resolution, by estimating the intrinsic uncertainty from cosmic variance (i.e. equation 9). After determining the targeted modes and frequencies, the second step is to predict the uncertainty introduced by the experimental setup (equation 16). The advantage of radio interferometry is that since each baseline is most sensitive to the fluctuation mode at a particular scale, we can choose a particular combination of baselines to tailor the analysis. For example, if the frequency anticorrelation parameter $\xi$ for a theory is highly dependent on $\ell$, so we want to constrain this theoretical model only in a particular smaller range of $\ell$ where $\xi_{\ell}\simeq\xi^{\prime}$ is approximately a constant, then we can down-weight the response to unwanted $\ell$ modes by filtering the baseline.
@@
In the cosmic variance-limited (noise-free) case where all of the spherical harmonic modes are measured within a band between $20\leq\ell\leq 90$ (or $40\leq\ell\leq 70$; see Table 1), we found that all of the foreground component parameters could be strongly constrained except for the (faint) Extragalactic Free-Free emission (Table 2). This is in the case where there is no fiducial RSB excess model present, and no prior has been applied. This suggests that close-packed arrays of the kind used for 21cm surveys can usefully disentangle different anisotropic components of the radio sky based on angular and spectral clustering alone, i.e. without absolute spectrometry.
@@
We also showed predictions for a HERA-like array to measure the angular power spectrum of the RSB excess for different toy models (results in Sec. 3.3). If the fiducial amplitude of the RSB excess component is of the same order as the dominant Galactic synchrotron component or larger, its amplitude and spectral/spatial power law indices can be constrained reasonably well within a typical observing season of $\sim 40$ nights or so, even if the spectral/spatial clustering properties are very similar to the Galactic synchrotron’s. This is only the case if $\sim 10\%$ priors are applied to the non-RSB component parameters, as otherwise significant degeneracies arise. If the RSB excess is fainter than this at the target frequencies, however, it becomes difficult to disentangle from the foreground components based on angular and spectral clustering properties alone, and so either substantially more observing time, or stronger priors on the other components would be needed to permit its characterisation in terms of the parameters of our assumed APS model.